OpenStudy (anonymous):

2(4x-3)-8=4+2x Please show to work this problem.

OpenStudy (anonymous):

When I tried to work and solve this equation I got: 2(4x-3)-8=4+2x 8x-6-8=4+2x and I got stuck there and now cannot solve nor show anymore work for the problem.

OpenStudy (anonymous):

is anyone willing to help me on this one?

OpenStudy (anonymous):

Yall act like my teacher, no help when i really need it